A restaurant donated 6 identical gift cards to be raffled off at a local charity event. At the charity's fundraising event, 12 people buy a raffle ticket to try to win one of the gift cards, and their raffle tickets are placed into a single can. The raffle organizer then randomly picks 6 tickets from the can to determine the gift card winners. How many possible outcomes are there?


Found 2 solutions by xxLaurax, Jjona5:
Answer by xxLaurax(5)

Remember permutation is when order does matter. So 12 people bought the raffle ticket and ONLY 6 people will get pick.
So, first person who gets pick can be anyone of the twelve people.
Second person who gets pick can be anyone of the eleven people.
And this goes on and on once you pick out the sixth person.
P = 12*11*10*9*8*7 = 665,280
You're welcome.

Answer by Jjona5(1)


So wouldnt it be 12!/(6!(12-6)!)=924 It should be a combination not a permutation?